OpenStudy (anonymous):

How do you raise a complicated fraction like 5x^4/4y^8 to the power of -2?

OpenStudy (amistre64):

the rules dont change by complexity

OpenStudy (amistre64):

\[(B^m)^n=B^{m\ *\ n}\]

OpenStudy (anonymous):

But those aren't negative...

OpenStudy (amistre64):

they are "any" numbers

OpenStudy (anonymous):

So I just add negative two the exponents already present?

OpenStudy (amistre64):

\[(\frac{5\ x^4}{4\ y^8})^{-2}\] \[\frac{(5)^{-2}\ (x^4)^{-2}}{(4)^{-2}\ (y^8)^{-2}}\] \[\frac{5^{-2}\ x^{4(-2)}}{4^{-2}\ y^{8(-2)}}\] \[\frac{5^{-2}\ x^{-8}}{4^{-2}\ y^{-16}}\] and flip as needed

OpenStudy (anonymous):

Okay, times them - and 'flip as needed', flip to get the exponents to be positive?

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

Okay! Thank you!

OpenStudy (amistre64):

youre welcome